OpenStudy (anonymous):

Which identity is not used in the proof of the identity? 1 + cot^2 theta = csc^2 theta

OpenStudy (anonymous):

here is the proof: http://puu.sh/3cMf2.png

OpenStudy (campbell_st):

a simple proof is to start with \[\sin^2(x)+ \cos^2(x) = 1\] divide each term by sin^2(x) gives \[1 + \frac{\cos^2(x)}{\sin^2(x)} = \frac{1}{\sin^2(x)}\] which gives \[1 + \cot^2(x) = \csc^2(x)\] so 1 ratio not used is tan^2(x) and perhaps sec^2(x)

OpenStudy (anonymous):

i need to know which identity is not used to find the proof. for example, cotangent identity, Pythagorean identity, etc

OpenStudy (campbell_st):

well the proof doesn't use sec(x) or tan(x) and it seems a complicated proof...

OpenStudy (anonymous):

my options are cotangent identity Pythagorean identity reciprocal identity and tangent identity

OpenStudy (anonymous):

TAngent Identity. It means the usage of tan

OpenStudy (anonymous):

Cotangent means the using of various properties of cot: cot^2x+1=cosec^2x Pythagoran is sin^2x+cos^2x=1 Reciprocal is 1/sin^2x=csc^2x So only tangent is left. :)

OpenStudy (anonymous):

I love medals! xD

OpenStudy (anonymous):

can you help me with one more?

OpenStudy (anonymous):

http://puu.sh/3cNKR.png

OpenStudy (anonymous):

@kutabs

OpenStudy (anonymous):

Option 3 is the only correct one the rest are wrong. Eg: sin^2x=cos^2x-1 which gives, cos^2x-sin^2=1 (Not possible as sin^2x+cos^2x=1 Phthagoran identity). Similarly you could prove the others wrong.

OpenStudy (anonymous):

Thank you so much!